OpenStudy (anonymous):

Which of the following is the equation of a line parallel to the line y = 4x + 1, passing through the point (5,1)? A. 4x + y = 19 B. 4x - y = 19 C. 4x + y = -19 D. -4x - y = 19

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

\[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

pooja195 (pooja195):

\[\huge~\rm~y-1=4(x-5)\] Is what you get can you make this equation into slope intercept form?

OpenStudy (anonymous):

New line must have a slope m= 4 (since parallel) and pass thru the points (5,1) new line y = mx + b y = 4x + b 1 = 4*5 + b 1 = 20 + b b= -19 new line is y = 4x - 19 rearrange 4x - y = 19

pooja195 (pooja195):

Unless the problem has specified what form it wants it in. @peak Please only post relevant comments..

OpenStudy (anonymous):

do you know lionel cruz selena929

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

B?

pooja195 (pooja195):

correct

OpenStudy (anonymous):

B?